OpenStudy (crazysuperanya):

Simplify: 2x + 5(x – 2) = 3(x – 5) PLEASE HELP

Mehek (mehek14):

first use distributive property for \(5(x-2)\)

Mehek (mehek14):

\(5*x= \\ 5*2=\)

Mehek (mehek14):

^answer that for me

OpenStudy (crazysuperanya):

hold on

OpenStudy (crazysuperanya):

10

Mehek (mehek14):

ok 5 * 2 = 10 what about 5*x ?

OpenStudy (crazysuperanya):

um... that depends what x is

Mehek (mehek14):

\(5*x = 5x\)

OpenStudy (crazysuperanya):

oh ok

Mehek (mehek14):

so you have \(2x + 5x - 10 = 3(x-5)\)

Mehek (mehek14):

what is \(3*x=\\3*(-5)\)

OpenStudy (crazysuperanya):

15

Mehek (mehek14):

-15

OpenStudy (crazysuperanya):

right sorry

Mehek (mehek14):

and 3*x

OpenStudy (crazysuperanya):

3*X=3X

Mehek (mehek14):

ok so now you have \(2x+5x-10=3x-15\) what is \(2x+5x\)

OpenStudy (crazysuperanya):

7x...?

Mehek (mehek14):

\(7x-10=3x-15\) \(7x-3x=?\)

OpenStudy (crazysuperanya):

4x

Mehek (mehek14):

\(4x-10=-15\) \(-15+10=?\)

OpenStudy (crazysuperanya):

-5

Mehek (mehek14):

\(\dfrac{-5}{4}\)

OpenStudy (crazysuperanya):

1.25

Mehek (mehek14):

-1.25

Mehek (mehek14):

so \(x=-1.25\)

OpenStudy (crazysuperanya):

Thank you so much i never would have gotten this without your help

Mehek (mehek14):

yw \(\ddot\smile\)

OpenStudy (crazysuperanya):

Could you help me with one more

Mehek (mehek14):

sorry but I gtg

OpenStudy (crazysuperanya):

oh... ok i fanned you though so maybe you'll be able to help another time!

Mehek (mehek14):

sure just tag me when I come on

OpenStudy (crazysuperanya):

ok again thank you for your help!!!!!!!!!!!